Angles 1 and 2 are a linear pair, the measure of angle 1=x-39, and the measure of angle 2=x+61. Find the measure of each angle.
Answer by stanbon(75887)

Angles 1 and 2 are a linear pair, the measure of angle 1=x-39, and the measure of angle 2=x+61. Find the measure of each angle.
------
Equation:
x-39 + x+61 = 180
2x + 22 = 180
x + 11 = 90
x = 79
---
Ans: x-39 = 40 degrees
x+61 = 140 degrees
============================
Cheers,
Stan H.
================